# Galaxy assembly bias in cosmological hydrodynamical simulations – a comparison between SIMBA and IllustrisTNG

### v1

@@ 7 Summary and Conclusions
Our results suggest that the dominant contributing factor to GAB is expected to be the enhanced likelihood for low-mass haloes in high-density regions to host galaxies. In comparison, we find that modifications of galaxy populations by the local tidal anisotropy are expected to be sub-dominant to the effect of overdensity. This prediction is in conflict with some claims based on observations (Alam et al., 2023); further studies of this discrepancy will be important in order to establish with certainty the empirical degree of assembly bias, and to gain a better understanding of its origin.
@@ Acknowledgements
HGY is grateful for support from the China Scholarship Council (CSC) under Grant No. CSC 202306340042. We are grateful to Shadab Alam and to Sergio Contreras for helpful comments that allowed us to tighten some of our arguments.

### v2

@@ 7 Summary and Conclusions
Our results suggest that the dominant contributing factor to GAB is expected to be the enhanced likelihood for low-mass haloes in high-density regions to host galaxies. Taken together with CAMELS results on the differing baryonic physics of SIMBA and TNG, our findings suggest that low-mass haloes in high-overdensity environments are more likely to form galaxies; however, this tendency can be suppressed by different recipes for AGN feedback from nearby massive haloes. In comparison, we find that modifications of galaxy populations by the local tidal anisotropy are expected to be sub-dominant to the effect of overdensity. This prediction is in conflict with some claims based on observations (Alam et al., 2023); further studies of this discrepancy will be important in order to establish with certainty the empirical degree of assembly bias, and to gain a better understanding of its origin.
@@ Acknowledgements
HGY is grateful for support from the China Scholarship Council (CSC) under Grant No. CSC 202306340042. We are grateful to Shadab Alam and to Sergio Contreras for helpful comments that allowed us to tighten some of our arguments. We thank the anonymous referee for their constructive comments, which helped us improve the discussion of our results.